OpenStudy (anonymous):

A restaurant earns $1073 on Friday and $1108 on Saturday. Write and solve an equation to find the amount x (in dollars) the restaurant needs to earn on Sunday to average $1000 per day over the three-day period. Write your equation so that the units on each side of the equation are dollars per day.

OpenStudy (australopithecus):

The question is asking about averages, The formula to calculate average is: \[Average = \frac{m_1 + m_2 +...}{n}\] Where n = the total number of events (in this case that would be the number of days) m = value of a single event Example: so if you made 1 dollars on day 1 and 3 dollars on day 2, the average money you make will be, n = 2 (since there are two days), m_1 = $1 and m_3 = $3 Average made between the two days = (1 + 3)/2 = $2 per day average In your case you have the average and you have the values for two days, but you dont have the value for the third day. So set up the average equation and solve for the unknown.

OpenStudy (anonymous):

Thank you! This helped a lot!